A corporation has a beta of 2. The firm expects to pay a dividend of $3 next year. This dividend is expected to continue to grow indefinitely at a constant rate of 3% per year. The risk free rate is 5%. The market portfolio has an expected return of 14%.

a) Calculate the required return on the corporation's shares.

b) Calculate the corporation's share price today.
Answer #1

a.Required return=risk free rate+beta*(market rate-risk free rate)

=5+2*(14-5)

=23%

b.Current price=D1/(Required return-Growth rate)

=3/(0.23-0.03)

=$15